# CleanMel: Mel-Spectrogram Enhancement for Improving Both Speech Quality and ASR

### v1

@@ V-C Influence of neural vocoder to ASR performance
We use a GAN-based neural vocoder to transform the enhanced Mel-spectrogram to waveform, which may modify the original Mel-spectrogram and reduce the fidelity, although a Mel-spectrogram loss is adopted in the vocoder to maintain the Mel-spectrogram fidelity. Table VII compares the ASR results between taking as input the enhanced Mel-spectrogram and the vocoder waveform. Note that the ASR models used for VoiceFixer is re-trained accroding to its STFT and Mel-frequency setups. We can see that the neural vocoders indeed reduce the Mel-spectrogram fidelity and thus the ASR performance. However, the performance reductions are not significant, and the ASR results of vocoder waveform for each model closely insist on the results of its corresponding enhanced Mel-spectrogram. This means the vocoder waveform could be also an alternative/suboptimal choice for ASR. If it is not easy to re-train a CleanMel model according to the setups of a new ASR model, since vocoder waveform is irrelevant to those setups, our pre-trained checkpoints can be used.
V-D Model size and computational complexity
TABLE VIII: Model size and computational complexity of different models. For CleanMel and VoiceFixer, the #Param and FLOPs of Mel-spectrogram enhancement network and vocoder are respectively given in addition, presenting the computational cost for ASR (only the former term) and speech enhancement (the addition).
Mode Model #Param(M) FLOPs(G/s)
Online FullSubNet [15] 5.6 59.6
Demucs [50] 33.5 15.4
oSpatialNet [22] 1.7 36.8
CleanMel-S (ours) 2.7 + 13.2 18.1 + 1.7
Offline FullSubNet [15] 14.6 157.9
VoiceFixer [19] 88.3 + 33.8 20.9 + 103.0
@@ V-D Model size and computational complexity
SpatialNet [21] 1.6 46.3
SpatialNet-Mamba 1.7 34.9
CleanMel-S (ours) 2.5 + 13.2 32.9 + 3.3
CleanMel-L (ours) 7.2 + 13.2 127.8 + 3.3
Table VIII shows the model size and computational complexity of the proposed model and comparison models. The proposed model is composed of a Mel-spectrogram enhancement network and an optional neural vocoder. The Mel-spectrogram enhancement network has a small model size and a large computational complexity, mainly due to the independent computation of frequencies with shared narrow-band blocks. Compared to linear-frequency sub-band/narrow-band processing networks, including FullSubNet, oSpatialNet, SpatialNet(-Mamba), the proposed Mel-spectrogram enhancement network has a smaller computational complexity, due to the much less frequencies to be processed, i.e. 80 Mel frequencies versus 256 linear frequencies. As a diffusion model, the computation of StoRM is much more expensive than other models.

### v2

@@ V-E Influence of neural vocoder to ASR performance
We use a GAN-based neural vocoder to transform the enhanced Mel-spectrogram to waveform, which may modify the original Mel-spectrogram and reduce the fidelity, although a Mel-spectrogram loss is adopted in the vocoder to maintain the Mel-spectrogram fidelity. Table VIII compares the ASR results between taking as input the enhanced Mel-spectrogram and the vocoder waveform. Note that the ASR models used for VoiceFixer is re-trained according to its STFT and Mel-frequency setups. We can see that the neural vocoders indeed reduce the Mel-spectrogram fidelity as well as the ASR performance. However, the performance reductions are not significant, and the ASR results of vocoder waveform for each model closely align with the results of its corresponding enhanced Mel-spectrogram. This means the vocoder waveform could be also an alternative/suboptimal choice for ASR. If it is not easy to re-train a CleanMel model according to the setups of a new ASR model, since vocoder waveform is irrelevant to those setups, our pre-trained checkpoints can be used.
TABLE IX: ASR results (WER, %) on CHiME4 of fine-tuning CleanMel and/or the ASR model.
Enhancement
method
Fine-tuning CHiME4
CleanMel ASR simu. real
unprocessed - 15.3 13.1
CleanMel-S-mask $\times$ $\times$ 11.9 9.3
$\times$ 10.6 8.8
$\times$ 11.2 8.8
11.1 8.8
TABLE X: Model size and computational complexity of different models.
Model #Param(M) FLOPs(G/s)
Online
FullSubNet [15] 5.6 60
Demucs [2] 33.5 15
LiSenNet [55] 0.04 0.1
oSpatialNet [22] 1.7 37
CleanMel-S + Vocos 2.7 + 13.2 18 + 2
Offline
ConvTasNet [1] 5.0 10
FullSubNet [15] 14.6 158
CMGAN [54] 1.8 62
VoiceFixer + HifiGAN [19] 88.3 + 33.8 21 + 103
CDiffuse [56] 10.2 1520
@@
UNIVERSE++ [58] 148.8 163
SpatialNet [21] 1.6 46
SpatialNet-Mamba 1.7 35
CleanMel-S + Vocos 2.5 + 13.2 33 + 3
CleanMel-L + Vocos 7.2 + 13.2 128 + 3
V-F Joint fine-tuning of the CleanMel and ASR models
As shown in the previous works [62, 63], joint fine-tuning of the cascaded speech enhancement and ASR models can mitigate speech distortions introduced by the speech enhancement model, and further improve the ASR performance. In this section, we investigate the fine-tuning strategies of the proposed model. Experiments are conducted on the CHiME4 dataset, with the pretrained CleanMel-S-mask model, and the pretrained E-branchformer ASR model of CHiME4.
Three strategies are evaluated: 1) freezing the ASR model and fine-tuning CleanMel-S-mask, 2) freezing CleanMel-S-mask and fine-tuning the ASR model and 3) joint fine-tuning of the CleanMel and ASR models. Fine-tuning is implemented in the ESPNet framework. We replace the STFT and logMel transformations of the original ASR system by our pretrained CleanMel model. Data augmentation of SpecAug [64] is applied to the enhanced logMel, but time-warping is not applied. All models are trained with the AdamW [51] optimizer. We observed that faster convergence is achieved when the ASR model is unfrozen. Accordingly, ASR-only fine-tuning and joint fine-tuning last for 2 epochs and CleanMel-only fine-tuning lasts for 4 epochs. For all experiments, the learning rate ramps up from 0 to 1e-5 during the first epoch (2,300 steps), then decays according to the principle set in ESPNet. Models are validated on the development set for every 230 steps. The best 10 checkpoints are averaged as the final model.
Results are presented in Table IX. It can be seen that ASR performance can be further improved by all strategies. This demonstrates that, although the pretrained CleanMel model can largely improve ASR performance over the unprocessed signal, the enhanced logMel still contains certain artifacts detrimental to ASR. These artifacts can be further mitigated through fine-tuning. Among the three strategies, only fine-tuning the CleanMel model achieves the best performance. However, this may not be conclusive, different results could be obtained for other ASR models or datasets.
V-G Model size and computational complexity
Table X shows the model size and computational complexity of the proposed model and comparison models. The proposed model is composed of a Mel-spectrogram enhancement network and an optional neural vocoder. The Mel-spectrogram enhancement network has a small model size and a large computational complexity, mainly due to the independent computation of frequencies with shared narrow-band blocks. Compared to linear-frequency sub-band/narrow-band processing networks, including FullSubNet, oSpatialNet, SpatialNet(-Mamba), the proposed Mel-spectrogram enhancement network has a smaller computational complexity, due to the much less frequencies to be processed, i.e. 80 Mel frequencies versus 256 linear frequencies. As for diffusion models, the computation cost is much more expensive than other models.